OpenStudy (anonymous):

please help. find the next three term sequence of this 32.5, 35.1, 37.7 _______, ______, _____.

OpenStudy (anonymous):

Do you know what the difference is?

OpenStudy (anonymous):

This is actually a very simple question when you think about it. the difference between the first two terms is 2.6, the difference between the 2nd and 3rd term is 2.6, because of the question we should assume that the function is linear because it doesn't give us any other data so the next term would be 37.7 + 2.6 = 40.3, the one after would be 40.3 + 2.6 = 42.9, and the last would be 42.9 + 2.6 = 45.5